# Rapid Dissipative Ground State Preparation at Chemical Transition States

### v1

@@ Abstract
Simulating chemical reactions is a central challenge in computational chemistry, characterized by an uneven difficulty profile: while equilibrium reactant and product geometries are often classically tractable, intermediate transition states frequently exhibit strong correlation that defies standard approximations. We present a protocol for dissipative ground state preparation that exploits this structure by treating the reaction path itself as a computational primitive. Our protocol uses an approach where a state prepared at a tractable geometry is propagated along a discretized reaction coordinate using Procrustes-aligned orbital rotations and stabilized by engineered dissipative cooling. We show that for reaction paths satisfying a localized Eigenstate Thermalization Hypothesis (ETH) drift condition in the strongly correlated regime, the algorithm prepares ground states of chemical systems with $N_{o}$ orbitals to an energy error $\epsilon_{E}$ with a total gate complexity scaling as $\widetilde{O}(N_{o}^{3}/\epsilon_{E})$. We provide logical resource estimates for benchmark systems including FeMoco, Cytochrome P450, and Ru-based carbon capture catalysts.
@@ I Introduction
Figure 1: (a) The transition digraph associated with the Lindbladian-induced Markov process $P$ depicting the longest canonical path between an excited state and the ground state; this path length indicates that the Markov process will mix rapidly. The underlying transition graph is a numerical example of H4 molecule at its multi-reference square geometry (see Figs. 3 and 2); the transition graph has the same number of nodes as the FCI dimension $D=\binom{2N_{o}}{N_{e}}=70$ for a $\text{CAS}(4,4)$ active space in the STO-3G orbtial basis. The values above each arrow are the transition probabilities from $|E_{j}\rangle\to|E_{k}\rangle$, $P_{j\to k}$ along the longest path; in this case the longest path is $|E_{D-1}\rangle\to|E_{0}\rangle$ with length $\ell=4$ whose transition probabilities are bounded below by $P_{j\to k}\geq p_{\min}=0.01$. (b) An example of a Ru-based carbon-capture catalyst reaction VIII $\to$ VIII–IX (transition state) $\to$ IX of the multi-step reaction mechanism for CO2 reduction [1]; the 2D potential energy surface (PES) is representative of a more complex, multi-dimensional mechanism underlying the reaction. (c) The logical quantum circuit for implementing the purely dissipative map $e^{t\mathcal{L}}$ that sends the approximate ground state $\rho_{i}$ and geometry $\mathbf{R}(s_{i})$ to the next ground state along the reaction path $\rho_{i+1}$ where circuit $U_{H}$ denotes the block encoding of a chemical Hamiltonian which is used to implement the logical circuit described in Appendix H.
@@
In this work, we present dissipative evolution, a quantum algorithm for preparing ground states at strongly correlated TS geometries along reaction paths. At a high level, the algorithm approximates the evolution along a reaction path from a system with an easy-to-prepare ground state to a system of interest by applying dissipative cooling primitives at a sequence of successive geometries along the discretized reaction path. The flexibility of the approach lies in the choice of reaction path and in the design of the cooling primitive. For example, one may consider simple linear paths passing through the TS, as in traditional adiabatic simulation [2] or chemically motivated choices such as the minimum-energy path (MEP) on the potential energy surface [3]. In this work, we focus on preparing the TS ground states along the MEP, as these geometries are of primary interest in quantum chemistry; however, our algorithm will apply to any path which is smooth.
@@
We show that, provided the Hamiltonian and its ground state vary smoothly along the reaction path—a property we quantify in terms of the smoothness of the pathway—ground states at transition-state geometries can be prepared with only polynomial overhead relative to the cost of preparing ground states at equilibrium reactant or product geometries. Specifically, in addition to the cost of state preparation, our algorithm has a runtime complexity bounded above by $\widetilde{O}(N_{o}^{3}/\epsilon_{E})$ where $\epsilon_{E}$ is the final energy error and $N_{o}$ is the number of (spin) orbitals. In practice, the complexity of our algorithm depends highly on the smoothness of the Hamiltonian along reaction paths which can be quantified with a Davis-Kahan-like constant $C_{\text{DK}}$. Physically, this constant quantified how quickly the ground state “rotated” along the reaction path. Our algorithm’s complexity in these terms becomes $\widetilde{O}(C_{\text{DK}}^{2}N_{o}/\epsilon_{E})$. Therefore, our algorithm obtains a substantial speed-up over the worst-case complexity bound precisely when the reaction path admits a smoothly varying Hamiltonian, and thereby, a smoothly evolving electronic ground state which is typically the case within the Born-Oppenheimer regime.
@@
Tensor network methods and quantum algorithms for ground state preparation have now matured to the point at which we can obtain accurate ground states at equilibrium geometries [4, 5]. Our algorithm amplifies the utility of this capability to prepare ground states at reactants and products, given that the real difficulty in accurately modeling quantum chemical systems is obtaining accurate TS ground states, and thereby accurate energetics, during bond breaking/bond forming. Our algorithm allows quantum computers to probe bond breaking/bond forming which is the very essence of chemistry, thus opening up an entirely new class of chemical problems with provable quantum advantage. This provides further evidence supporting an affirmative answer to the long-standing question of whether verifiable exponential quantum advantage exists in quantum chemistry for certain regimes [6].
@@
Quantum computers offer the potential to provide high-accuracy approximations to ground states even in these strongly-correlated regimes through algorithms such as quantum phase estimation (QPE) [7, 8, 9, 10, 11, 12, 13, 14, 15] and quantum imaginary time evolution (QITE) [16]. Yet, the performance of these approaches is constrained by the availability of a good initial guess, a state with nontrivial overlap with the target ground state  [17, 18, 19]. Recent progress on engineered open-system (Lindbladian) protocols has also enabled thermal and ground state preparation for chemical Hamiltonians [20, 21]. In particular [20] provides detailed analysis on cooling sub-routines that we build on here by applying to transition state chemistry and modifying the family of filter functions used. These methods, too, require the provision of a “warm start,” an initial guess that has some guarantee to be close to the true ground state in order to avoid sub-optimal local minima in the energetic landscape [22]. Providing such an approximation is far from trivial, as the electronic-structure Hamiltonian is known to be QMA-complete and NP-Hard depending on the setting [23, 24, 25]. This worst-case hardness motivates the question of whether chemically structured instances admit defensible quantum speedups in practice [6], especially those with clear scientific or economic value.
@@
The standard quantum algorithmic approach is to determine the reaction path classically using density functional theory (DFT) calculations, applying QPE to an ansatz state that can be efficiently described classically at a fixed geometry along the path. [26]. The success of this approach relies on an initial state with significant overlap with the true ground state; however, in strongly correlated regimes the overlap can be exponentially small in the number of system qubits, rendering QPE inefficient as a ground-state preparation routine on its own. This motivates end-to-end strategies that propagate a warm start along a reaction path, maintaining efficiency as the system enters the strongly correlated transition-state region separating reactants and products.
Very recently, Puig et al. [27] proposed a similar warm-start strategy that tracks a discretized Hamiltonian deformation along an arbitrary path through configuration space. The approach of Puig et al. [27] uses a variational ground-state preparation technique at each Hamiltonian along the discritized path, using each solution to initialize the next. The authors argue that this can improve trainability when the path avoids gap closings. Our work shares the same path-tracking intuition, but targets the transition-state bottleneck with an engineered dissipative cooling primitive designed to stabilize a transported warm start as correlations strengthen.
@@
Our work complements recent progress on dissipative and hybrid strategies for chemical state preparation along reaction coordinates, including constructions that improve warm starts and enable controlled access to excited manifolds [28, 29]. In this emerging picture, equilibrium-quality classical (or hybrid) preparation supplies high-fidelity endpoints, while path-based dissipative evolution provides a principled mechanism to carry those warm starts through the transition-state bottleneck where classical guiding states are least reliable.
The remainder of the paper is organized as follows. In Section II, we prove the efficiency i.e., identify the regime in which our algorithm requires $O(\mathrm{poly}(N_{o}))$ quantum resources. In Section III provide quantum resource estimates associated with running our algorithm on real-world quantum chemical problems on fault tolerant quantum computers. In Section IV, we discuss how our algorithm fits into the on-going pursuit of developing practical fault tolerant quantum algorithms for studying chemical reactions. In the Appendices B and F, we provide proofs of the theorems introduced herein, in addition to the methodology behind our resource estimation for practical quantum circuit implementations of our algorithm in Appendices I and H. Additionally, we establish our algorithm within the broader landscape of ground-state preparation methods through a detailed comparative analysis in Appendices C and D.
@@ II.1 Definitions and Assumptions
where $M$ scales as $O(N_{o})$ for electronic structure Hamiltonian and $N_{o}$ is the number of (spatial) orbitals in the system [30]. In this frame, we use a Lipschitz smoothness criteria:
@@
The Davis-Kahan constant $C_{\text{DK}}$ is analogous to the “adiabatic condition” that appears in adiabatic ground state preparation [31, 2], where the ratio of the path derivative to the instantaneous ground state gap controls how rapidly the ground state can change with $s$.
@@
which partitions the spectrum into layers of equal energy thickness. We say that $\mu$ is $\Delta E_{\text{layer}}$-concentrated if there exist $L_{\mu}\in\mathbb{N}$ and $\eta\in(0,1)$ such that
@@
With these definitions, the population dynamics induced by a dissipative primitive can be reduced to a classical Markov chain $(J_{t})_{t\geq 0}$ on the eigenstate labels. Following [20, 32], the transition kernel is a row-stochastic matrix $P$ mapping the pre-primitive distribution $\mu$ to the post-primitive distribution $\mu^{\prime}$, i.e., $\mu^{\prime}=\mu P$. For a formal definition, see Appendix A. The matrix $P$ can also be viewed as an edge-weighted directed graph (see Fig. 1a).
@@ II.2 Efficient Ground State Preparation at Transition Geometries
In this section, we demonstrate how dissipative evolution algorithm allows for efficient quantum state preparation, using open quantum system simulation as an algorithmic primitive [33, 20, 21]. In a chemical reaction, the ground states of Hamiltonians at the reactant or product geometries are typically well described by classical means. At the transition-state geometry, however, $H(s^{*})$ is typically too strongly interacting for classical methods to provide any guarantees. Thus, our goal is to design a quantum algorithm that prepares a state $\rho$ such that
@@
At the algorithmic level, we proceed by continuation along the reaction path. We discretize $s\in[0,1]$ into a sequence of geometries and, at each step, apply a dissipative cooling primitive that amplifies the ground-state population at the new geometry, using the ground-state approximation from the previous step as input (see Fig. 1b). The overall complexity is governed primarily by how rapidly the ground-state subspace varies with $s$ and by the target energy accuracy $\epsilon_{E}$.
@@
Ground-state preparation at transition-state geometries is known to be a computationally challenging problem for which a myriad of quantum algorithms have been designed. Notable examples include quantum phase estimation [7] (QPE), digital adiabatic simulation [2] (DAS), phase randomization [31] (PR) and dynamic cooling [34] (DC). Table 1 presents a comparison between the asymptotic runtime of these algorithms with dissipative evolution in terms of their shared parameters: the number of orbitals $N_{o}$, the maximum Hamiltonian norm along the reaction path $\|H\|\coloneqq\max_{s}\|H(s)\|$, and the minimum spectral gap $\Delta_{\min}\coloneqq\min_{s}\Delta(s)$. In terms of these parameters, dissipative evolution, using a Lindbladian-type cooling primitive has a gate complexity of
@@
QPE [7] $\widetilde{O}\!\left(\frac{\|H\|}{\Delta_{\min}}\cdot\frac{1}{\epsilon_{E}}\cdot\frac{1}{p_{0}}\cdot\mathrm{poly}(N_{o})\right)$
DAS [2] $\widetilde{O}\!\left(\frac{\|H\|^{3}}{\Delta_{\min}^{5}}\cdot\frac{1}{\epsilon_{E}}\cdot\mathrm{poly}(N_{o})\right)$
PR [31] $\widetilde{O}\!\left(\frac{\|H\|}{\Delta_{\min}^{3}}\cdot\frac{1}{\epsilon_{E}}\cdot\frac{1}{1-p_{\mathrm{succ}}}\cdot\mathrm{poly}(N_{o})\right)$
DC [34] $\widetilde{O}\!\left(\frac{\|H\|}{\Delta_{\min}^{3}}\cdot\frac{1}{\epsilon_{E}}\cdot\mathrm{poly}(N_{o})\right)$
@@
Table 1: Gate complexity comparison (suppressing polylogarithmic factors via $\widetilde{O}$). Here $\|H\|$ denotes that maximum operator norm over the reaction path, and $\Delta_{\min}$ is the minimum spectral gap along the path. The target energy precision is $\epsilon_{E}$. $N_{o}$ is a problem-size parameter (e.g., number of spin orbitals), and $\mathrm{poly}(N_{o})$ denotes method-dependent polynomial overheads (such as basis changes, state preparation subroutines, or term-selection costs). For QPE, $p_{0}\coloneqq|\langle E_{0}(s^{*})|\phi\rangle|^{2}$ is the initial squared overlap of the prepared state $|\phi\rangle$ with the target ground state $|E_{0}(s^{*})\rangle$ at the hardest point $s^{*}$, giving the expected $1/p_{0}$ repetition overhead. For PR, $p_{\mathrm{succ}}$ is the per-attempt success probability of the projective/randomization step, yielding repetition overhead $1/(1-p_{\mathrm{succ}})$. For our method, we use the estimate $C_{\mathrm{DK}}=O(N_{o}/\Delta_{\min})$, which is absorbed into the displayed scaling.
@@
The asymptotic runtime complexity of dissipative evolution is quadratically better than DAS in terms of the Hamiltonian norm over the minimum gap $\|H\|/\Delta_{\text{min}}$. A key distinction between DAS and dissipative evolution is how errors scale and accumulate along the path. Digital adiabatic simulation is a coherent transport method: to keep the state within the instantaneous low-energy subspace, the schedule must suppress diabatic transitions, and existing worst-case bounds therefore exhibit a strong dependence on the minimum spectral gap along the path (e.g., the $\Delta_{\min}^{-5}$ dependence in [2]). By contrast, dissipative evolution implements a contractive channel that relaxes toward the instantaneous low-temperature state (and, in the zero-temperature limit, toward the ground-state manifold). In the regime we target—where a warm start is available at an easy geometry and the path is sufficiently smooth—state-preparation reduces to repeatedly re-cooling at intermediate geometries. Under the uniform downward drift / mixing assumptions used in our analysis, the number of cooling steps is controlled by a mixing-time parameter (captured in our constant $C_{\mathrm{DK}}$ and related hitting-time bounds) rather than by an adiabatic gap condition. Accordingly, our advantage over DAS should be understood as replacing worst-case adiabatic gap dependence with a mixing-time dependence that can be substantially milder in ETH-like, locally thermalizing regions; we do not claim an unconditional improvement in $\|H\|/\Delta_{\min}$ in the worst case.

### v2

@@ Abstract
Simulating chemical reactions exhibits a pronounced unevenness in computational difficulty: while equilibrium reactant and product geometries are often tractable, transition-state (TS) geometries frequently display strong multi-reference character that challenges both classical solvers and coherent quantum state-preparation methods. We introduce a dissipative continuation protocol for preparing electronic ground states along reaction paths that targets this transition-state bottleneck. Starting from a warm start at a tractable geometry, the algorithm transports the state along a discretized reaction coordinate using orbital-aligned Hamiltonians and engineered dissipative cooling primitives that repeatedly contract population into the instantaneous low-energy sector. We prove that, for reaction paths satisfying a Lipschitz smoothness condition and a localized downward-drift (ETH-motivated) mixing assumption within the relevant energy window, the ground state at a target geometry can be prepared to energy error $\epsilon_{E}$ with total gate complexity $\widetilde{O}(C_{DK}^{2}N_{o}/\epsilon_{E})$, where $C_{DK}$ quantifies ground-state rotation along the path. This identifies a structured chemical regime in which dissipative continuation provides a controlled and scalable route to ground-state preparation at strongly correlated transition-state geometries.
@@ I Introduction
Figure 1: (a) The transition digraph associated with the Lindbladian-induced Markov process $P$ depicting the longest canonical path between an excited state and the ground state; this path length indicates that the Markov process will mix rapidly. The underlying transition graph is a numerical example of H4 molecule at its multi-reference square geometry (see Figs. 4 and 3); the transition graph has the same number of nodes as the FCI dimension $D=\binom{2N_{o}}{N_{e}}=70$ for a $\text{CAS}(4,4)$ active space in the STO-3G orbtial basis. The values above each arrow are the transition probabilities from $|E_{j}\rangle\to|E_{k}\rangle$, $P_{j\to k}$ along the longest path; in this case the longest path is $|E_{D-1}\rangle\to|E_{0}\rangle$ with length $\ell=4$ whose transition probabilities are bounded below by $P_{j\to k}\geq 0.01$. (b) An example of a Ru-based carbon-capture catalyst reaction VIII $\to$ VIII–IX (transition state) $\to$ IX of the multi-step reaction mechanism for CO2 reduction [vonBurg2021]; the 2D potential energy surface (PES) is representative of a more complex, multi-dimensional mechanism underlying the reaction. (c) The logical quantum circuit for implementing the purely dissipative map $e^{t\mathcal{L}}$ that sends the approximate ground state $\rho_{i}$ and geometry $\mathbf{R}(s_{i})$ to the next ground state along the reaction path $\rho_{i+1}$ where circuit $U_{H}$ denotes the block encoding of a chemical Hamiltonian which is used to implement the logical circuit described in Appendix H.
@@
In this work, we present dissipative evolution, a quantum algorithm for preparing ground states at strongly correlated TS geometries along reaction paths. At a high level, the algorithm approximates the evolution along a reaction path from a system with an easy-to-prepare ground state to a system of interest by applying dissipative cooling primitives at a sequence of successive geometries along the discretized reaction path. The flexibility of the approach lies in the choice of reaction path, which can be modified and optimized for desirable properties and in the design of the cooling primitive. For example, one may consider simple linear paths passing through the TS, as in traditional adiabatic simulation [Berry2025Large] or chemically motivated choices such as the minimum-energy path (MEP) on the potential energy surface [Nguyen2026]. In this work, we focus on preparing the TS ground states along the MEP, as these geometries are of primary interest in quantum chemistry; however, our algorithm will apply to any path which is smooth.
@@
We show that, provided the Hamiltonian and its ground state vary smoothly along the reaction path—a property we quantify in terms of the smoothness of the pathway—ground states at transition-state geometries can be prepared with only polynomial overhead relative to the cost of preparing ground states at equilibrium reactant or product geometries. Specifically, in addition to the cost of state preparation, our algorithm has a runtime complexity bounded above by $\widetilde{O}(N_{o}^{3}/\epsilon_{E})$ where $\epsilon_{E}$ is the final energy error and $N_{o}$ is the number of (spatial/spin) orbitals. In practice, the complexity of our algorithm depends highly on the smoothness of the Hamiltonian along reaction paths which can be quantified with a Davis-Kahan-like constant $C_{\text{DK}}$. Physically, this constant quantifies how quickly the ground state “rotated” along the reaction path. Our algorithm’s complexity in these terms becomes $\widetilde{O}(C_{\text{DK}}^{2}N_{o}/\epsilon_{E})$. Therefore, our algorithm obtains a substantial speed-up over the worst-case complexity bound precisely when the reaction path admits a smoothly varying Hamiltonian, and thereby, a smoothly evolving electronic ground state which is typically the case within the Born-Oppenheimer regime.
While our protocol transports a state along a reaction path, it should not be viewed as a dissipative analogue of adiabatic continuation. Adiabatic approaches rely on coherent evolution and require suppressing diabatic transitions globally along the path, leading to worst-case runtime bounds governed by small spectral gaps. In contrast, our method performs a sequence of local dissipative stabilization steps that repeatedly re-concentrate population into the instantaneous low-energy manifold. The role of the reaction path is therefore not to define a slow physical evolution, but to ensure geometric smoothness so that each cooling stage begins from a warm start with bounded excess energy. Algorithmic efficiency arises from mixing-time properties of the engineered Lindbladian rather than from adiabatic gap protection. In particular, path optimization is used here as a computational design principle: by selecting geometries that avoid abrupt projector rotations or strong bottlenecks, the continuation problem is converted from a coherent transport task into a sequence of locally contracting dissipative updates. This distinction places the present work conceptually closer to continuation methods with stabilization than to digital adiabatic simulation, even though both frameworks exploit smooth Hamiltonian deformations.
@@
Tensor network methods and quantum algorithms for ground state preparation have now matured to the point at which we can obtain accurate ground states at equilibrium geometries [zhai2026classicalsolutionfemoco, Goings_2022]. Our algorithm amplifies the utility of this capability to prepare ground states at reactants and products, given that the real difficulty in accurately modeling quantum chemical systems is obtaining accurate TS ground states, and thereby accurate energetics, during bond breaking/bond forming. Our algorithm allows quantum computers to probe bond breaking/bond forming, which is the very essence of chemistry, thus opening up an entirely new class of chemical problems with provable quantum advantage. This work identifies a structured regime in which dissipative continuation yields polynomial overhead beyond equilibrium ground-state preparation [Lee_2023].
@@
Quantum computers offer the potential to provide high-accuracy approximations to ground states even in these strongly-correlated regimes through algorithms such as quantum phase estimation (QPE) [Kitaev1995AbelianStabilizer, ge2019faster, mande2023tight, AspuruGuzik2005MolecularEnergies, Whitfield2011ElectronicStructure, OMalley2016PRXQPE, Kassal2009PropertiesGeometry, Kassal2008ChemicalDynamics, Lanyon2010NatChemH2] and quantum imaginary time evolution (QITE) [Motta2020-qite]. Yet, the performance of these approaches is constrained by the availability of a good initial guess, a state with nontrivial overlap with the target ground state  [Cade2023, Gharibian2023, Waite2025Physical]. Recent progress on engineered open-system (Lindbladian) protocols has also enabled thermal and ground state preparation for chemical Hamiltonians [ding2024single, ding2025End]. In particular [ding2024single] provides detailed analysis on cooling sub-routines that we build on here by applying to transition state chemistry and modifying the family of filter functions used. These methods, too, require the provision of a “warm start,” an initial guess that has some guarantee to be close to the true ground state in order to avoid sub-optimal local minima in the energetic landscape [Chen2025]. Providing such an approximation is far from trivial, as the electronic-structure Hamiltonian is known to be QMA-complete and NP-Hard depending on the setting [OGorman2022Intractability, Schuch2009, Whitfield2013]. This worst-case hardness motivates the question of whether chemically structured instances admit defensible quantum speedups in practice [Lee_2023], especially those with clear scientific or economic value.
@@
The standard quantum algorithmic approach is to determine the reaction path classically using density functional theory (DFT) calculations, applying QPE to an ansatz state that can be efficiently described classically at a fixed geometry along the path. [Reiher_2017]. The success of this approach relies on an initial state with significant overlap with the true ground state; however, in strongly correlated regimes the overlap can be exponentially small in the number of system qubits, rendering QPE inefficient as a ground-state preparation routine on its own. This motivates end-to-end strategies that propagate a warm start along a reaction path, maintaining efficiency as the system enters the strongly correlated transition-state region separating reactants and products.
Very recently, puig2026warmstartscoldstates proposed a similar warm-start strategy that tracks a discretized Hamiltonian deformation along an arbitrary path through configuration space. The approach of puig2026warmstartscoldstates uses a variational ground-state preparation technique at each Hamiltonian along the discretized path, using each solution to initialize the next. The authors argue that this can improve trainability when the path avoids gap closings. Our work shares the same path-tracking intuition, but targets the transition-state bottleneck with an engineered dissipative cooling primitive designed to stabilize a transported warm start as correlations strengthen.
@@
Our work complements recent progress on dissipative and hybrid strategies for chemical state preparation along reaction coordinates, including constructions that improve warm starts and enable controlled access to excited manifolds [Li2025_excitedstates, Li_2025]. In this emerging picture, equilibrium-quality classical (or hybrid) preparation supplies high-fidelity endpoints, while path-based dissipative evolution provides a principled mechanism to carry those warm starts through the transition-state bottleneck where classical guiding states are least reliable.
The remainder of the paper is organized as follows. In Section II, we prove the efficiency i.e., identify the regime in which our algorithm requires $O(\mathrm{poly}(N_{o}))$ quantum resources, and show how to optimize the reaction path to ensure efficiency. In Section III we provide an implementation of our algorithm, we present a set of numerical models to validate the algorithm and provide evidence for the ubiquity of the ETH-like regime, and present resource estimates associated with running our algorithm on real-world chemical problems for fault-tolerant quantum computers. In Section IV, we discuss how our algorithm fits into the ongoing pursuit of developing practical fault tolerant quantum algorithms for studying chemical reactions. In Appendices B and F, we provide proofs of theorems introduced herein, in addition to the methodology behind our resource estimation for practical quantum circuit implementations of our algorithm in Appendices I and H. Additionally, we establish our algorithm within the broader landscape of ground-state preparation methods through a detailed comparative analysis in Appendices C and D.
@@ II.1 Definitions and Assumptions
where $M$ scales as $O(N_{o})$ for electronic structure Hamiltonian and $N_{o}$ is the number of (spatial) orbitals in the system [Motta_2021]. In this frame, we use a Lipschitz smoothness criteria:
@@
The Davis-Kahan constant $C_{\text{DK}}$ is analogous to the “adiabatic condition” that appears in adiabatic ground state preparation [Boixo2009Eigenpath, Berry2025Large], where the ratio of the path derivative to the instantaneous ground state gap controls how rapidly the ground state can change with $s$.
To optimize reaction paths, we must define an appropriate notion of path smoothness, this is captured by the operator-Lipschitz constant.
Definition 4 (Operator-Lipschitz constant).
Let $\mathbf{R}:[0,1]\to\mathbb{R}^{3N}$ be a $C^{1}$ reaction path with Fréchet derivative $D_{\mathbf{R}}H_{\mathrm{el}}(\mathbf{R}):\mathbb{R}^{3N}\to\mathsf{Herm}$. Define the operator-Lipschitz constant $L_{\max}$ as a scalar that bounds the operator norm of $D_{\mathbf{R}}H_{\mathrm{el}}(\mathbf{R})$,
$\sup_{\|v\|=1}\bigl\|D_{\mathbf{R}}H_{\mathrm{el}}(\mathbf{R})[v]\bigr\|_{2}\leq L_{\max}.$
@@
which partitions the spectrum into layers of equal energy thickness, Fig. 5 shows a depiction of the layers for increasing sizes of active spaces. We say that $\mu$ is $\Delta E_{\text{layer}}$-concentrated if there exist $L_{\mu}\in\mathbb{N}$ and $\eta\in(0,1)$ such that
@@
With these definitions, the population dynamics induced by a dissipative primitive can be reduced to a classical Markov chain $(J_{t})_{t\geq 0}$ on the eigenstate labels. Following [ding2024single, LevinPeresWilmer2009], the transition kernel is a row-stochastic matrix $P$ mapping the pre-primitive distribution $\mu$ to the post-primitive distribution $\mu^{\prime}$, i.e., $\mu^{\prime}=\mu P$. For a formal definition, see Appendix A. The matrix $P$ can also be viewed as an edge-weighted directed graph (see Fig. 1a).
@@ II.2 Efficient Ground State Preparation at Transition Geometries
In this section, we demonstrate how dissipative evolution algorithm allows for efficient quantum state preparation, using open quantum system simulation as an algorithmic primitive [kastoryano2016quantum, ding2024single, ding2025End]. In a chemical reaction, the ground states of Hamiltonians at the reactant or product geometries are typically well described by classical means. At the transition-state geometry, however, $H(s^{*})$ is typically too strongly interacting for classical methods to provide any guarantees. Thus, our goal is to design a quantum algorithm that prepares a state $\rho$ such that
@@
At the algorithmic level, we approximate the potential energy surface using classical methods and then optimize a reaction pathway that meets the smoothness condition. We then proceed by continuation along the reaction path. We discretize $s\in[0,1]$ into a sequence of geometries and, at each step, apply a dissipative cooling primitive that amplifies the ground-state population at the new geometry, using the ground-state approximation from the previous step as input (see Fig. 1b). The overall complexity is governed primarily by how rapidly the ground-state subspace varies with $s$ and by the target energy accuracy $\epsilon_{E}$.
Concretely, we first identify a chemical system of interest that comes equipped with one or more reaction pathways that smoothly connect points along the pathway(s). These reaction pathways are typically determined using a transition-state search algorithm such as nudged elastic band, synchronous transit-guided quasi-Newton, or constrained optimization [Williams2016]. Typically chemist’s look for the minimum energy pathway (MEP), however, we have the freedom to select any smooth pathway so long as it contains the physical transition state geometry. The opens up an entire design space for optimizing the runtime of our algorithm for a given chemical reaction of interest. We discuss the reaction pathway optimization via the calculus of variations in Section II.4.
Once a reaction path is selected and optimized, we discretize the path in to $N_{H}$ points $\{s_{i}\}_{i=0}^{N_{H}-1}$ where $s_{0}=0$ is the reactant/product equilibrium chemical geometry and $s_{N_{H}-1}=s^{*}$ is the target transition state chemical geometry. At each point along the reaction, we classically build electronic structure Hamiltonians $H_{\text{el}}(\mathbf{R}(s_{i}))$ then use one of many classical post-Hartree Fock methods, namely the density matrix renormalization group (DMRG), to generate effective warm start at the initial geometry $s_{0}$.
@@
Ground-state preparation at transition-state geometries is known to be a computationally challenging problem for which a myriad of quantum algorithms have been designed. Notable examples include quantum phase estimation [Kitaev1995AbelianStabilizer] (QPE), digital adiabatic simulation [Berry2025Large] (DAS), phase randomization [Boixo2009Eigenpath] (PR) and dynamic cooling [motlagh2024_cooling] (DC). Table 1 presents a comparison between the asymptotic runtime of these algorithms with dissipative evolution in terms of their shared parameters: the number of orbitals $N_{o}$, the maximum Hamiltonian norm along the reaction path $\|H\|\coloneqq\max_{s}\|H(s)\|$, and the minimum spectral gap $\Delta_{\min}\coloneqq\min_{s}\Delta(s)$. In terms of these parameters, dissipative evolution, using a Lindbladian-type cooling primitive has a gate complexity of
@@
QPE [Kitaev1995AbelianStabilizer] $\widetilde{O}\!\left(\frac{\|H\|}{\Delta_{\min}}\cdot\frac{1}{\epsilon_{E}}\cdot\frac{1}{p_{0}}\cdot\mathrm{poly}(N_{o})\right)$
DAS [Berry2025Large] $\widetilde{O}\!\left(\frac{\|H\|^{3}}{\Delta_{\min}^{5}}\cdot\frac{1}{\epsilon_{E}}\cdot\mathrm{poly}(N_{o})\right)$
PR [Boixo2009Eigenpath] $\widetilde{O}\!\left(\frac{\|H\|}{\Delta_{\min}^{3}}\cdot\frac{1}{\epsilon_{E}}\cdot\frac{1}{1-p_{\mathrm{succ}}}\cdot\mathrm{poly}(N_{o})\right)$
DC [motlagh2024_cooling] $\widetilde{O}\!\left(\frac{\|H\|}{\Delta_{\min}^{3}}\cdot\frac{1}{\epsilon_{E}}\cdot\mathrm{poly}(N_{o})\right)$
@@
Table 1: Gate complexity comparison (suppressing polylogarithmic factors via $\widetilde{O}$). Here $\|H\|$ denotes that maximum operator norm over the reaction path, and $\Delta_{\min}$ is the minimum spectral gap along the path. The target energy precision is $\epsilon_{E}$. $N_{o}$ is a problem-size parameter (e.g., number of spatial/spin orbitals), and $\mathrm{poly}(N_{o})$ denotes method-dependent polynomial overheads (such as basis changes, state preparation subroutines, or term-selection costs). For QPE, $p_{0}\coloneqq|\langle E_{0}(s^{*})|\phi\rangle|^{2}$ is the initial squared overlap of the prepared state $|\phi\rangle$ with the target ground state $|E_{0}(s^{*})\rangle$ at the hardest point $s^{*}$, giving the expected $1/p_{0}$ repetition overhead. For PR, $p_{\mathrm{succ}}$ is the per-attempt success probability of the projective/randomization step, yielding repetition overhead $1/(1-p_{\mathrm{succ}})$. For our method, we use the estimate $C_{\mathrm{DK}}=O(N_{o}/\Delta_{\min})$, which is absorbed into the displayed scaling.
@@
Dissipative evolution is intended for settings where constructing the guiding state $|\phi\rangle$ with large overlap with the ground-state is unavailable or unreliable. This situation commonly arises in strongly multi-reference regions [Wardzala2024, Szalay2012, Jiang2012, Sharma2021]. In these regimes the ground state is no longer well described by a single determinant or low-entanglement ansatz: several near-degenerate configurations contribute with comparable weight and natural-orbital occupations deviate significantly from idempotency as the nuclear geometry approaches a symmetry-enhanced or bond-rearrangement configuration. As a result, even chemically motivated ansätze prepared at TS geometries may have extremely poor overlap, making it challenging in practice to guarantee provably non-negligible overlap without extensive active-space optimization that may be exponentially expensive. In such cases, QPE’s cost may be dominated by the $1/p_{0}$ factor in Eq. 12.
@@
The asymptotic runtime complexity of dissipative evolution is quadratically better than DAS in terms of the Hamiltonian norm over the minimum gap $\|H\|/\Delta_{\text{min}}$. A key distinction between DAS and dissipative evolution is how errors scale and accumulate along the path. Digital adiabatic simulation is a coherent transport method: to keep the state within the instantaneous low-energy subspace, the schedule must suppress diabatic transitions, and existing worst-case bounds therefore exhibit a strong dependence on the minimum spectral gap along the path (e.g., the $\Delta_{\min}^{-5}$ dependence in [Berry2025Large]). By contrast, dissipative evolution implements a contractive channel that relaxes toward the instantaneous low-temperature state (and, in the zero-temperature limit, toward the ground-state manifold). In the regime we target—where a warm start is available at an easy geometry and the path is sufficiently smooth—state-preparation reduces to repeatedly re-cooling at intermediate geometries. Under the uniform downward drift / mixing assumptions used in our analysis, the number of cooling steps is controlled by a mixing-time parameter (captured in our constant $C_{\mathrm{DK}}$ and related hitting-time bounds) rather than by an adiabatic gap condition. Accordingly, our advantage over DAS should be understood as replacing worst-case adiabatic gap dependence with a mixing-time dependence that can be substantially milder in ETH-like, locally thermalizing regions; we do not claim an unconditional improvement in $\|H\|/\Delta_{\min}$ in the worst case.